OpenStudy (anonymous):

lim x^3 - (X^2/92x+1) X-->-(1/2) please help Ive been stuck on this question for along time

OpenStudy (anonymous):

this is what I interpreted: \[\lim_{x \rightarrow -0.5} x^3-{{x^2} \over {92x+1}}\] if it is this case, then just substitute x = -0.5 I got an answer -43/360

OpenStudy (anonymous):

sorry missed the typo its X^2/(2x+1)

OpenStudy (anonymous):

I can break it down to ((X^2)(2x-1)(x+1))/(2X+1) but then i get stuck

OpenStudy (valpey):

This limit diverges to negative infinity. You sure you are writing this right?

OpenStudy (valpey):

\[\lim_{x \rightarrow -1/2}x^3-\frac{x^2}{2x+1}?\]

OpenStudy (anonymous):

yes but it wouldnt suprise me that the teacher gave out a typo, he has done it many times before

OpenStudy (valpey):

Also, I should say that this function has a vertical asymptote at x = -1/2

OpenStudy (anonymous):

okay so then the limit is non existent then I should have seen that in my graph but looking at it now I miss typed it in there to thank you

OpenStudy (anonymous):

Ok first separate the easy things out \[\=lim_{x \rightarrow -0.5} x^3-\lim_{x \rightarrow -0.5} {{x^2} \over{2x+1}}\] Substitute the one term and apply long divison \[=-0.125-[{x \over 2}-{x \over {2(2x+1)}}]\] now the limit is pretty easy, divide top and bottom by x \[=-0.125-0.25 +\lim_{x \rightarrow -0.5} {x \over {4x+2}}\] \[=-0.375 +\lim_{x \rightarrow -0.5} {x \over {x(4+{2 \over x})}}\] now cancel and substitute \[=-0.375 + {1\over4+4} \] \[=-{1 \over 4}\]